• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name D.Wang

    Log entry time 21:15:56 on October 31, 2009

    Entry number 297066

    keyword=Left arm cosmic check running

    After reset the trigger supervisor, CODA runs fine. The current run is 25156 and we will keep it running over night.